OpenStudy (anonymous):

pls help me im in a hurry will give a medal -5/4n = 40 n=? there is no choice for this

OpenStudy (anonymous):

do u play rotmg u look like a famous yutubers friend

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-5/4n = 40 (-5/4n)(4n) = 40(4n) -5 = 160n -5/160 = 160n/160 -5/160 = n -1/32 = n

OpenStudy (anonymous):

thx